What is X= -Y
2Y-3x=25

Answer by mananth(16949)

1 x + 1 y = 0 .............1
-3 x + 2 y = 25 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-2 x -2 y = 0
-3 x 2 y = 25
Add the two equations
-5 x = 25
/ -5
x = -5
plug value of x in (1)
1 x + 1 y = 0
-5 + 1 y = 0
1 y = 0+5
1 y = 5
y = 5